A steel press tool with a mass of 15 kg and a temperature of 950
Answer #1

Let the final temperature be T.

Since the steel is in higher temperature heat will flow from steel to the oil.

Mass of Steel = Ms = 15 kg.

Temperature of steel = Ts = 950 C = 950+273 = 1233 K

Specific heat capacity of steel = Cs = 400 J/kgK

Volume of oil = 50 litres = 0.05 m^3

Mass of Oil = Mo = volume*density = 0.05 * 800 = 40 kg

Temperature of oil = To = 15 degree C = 15+273 = 288 K

Specific heat capacity of oil = Co = 1800 J/kgK

Heat flow from steel = heat gain by oil

MsCs(Ts-T) = MoCo(T-To)

15*400*(1233 - T) = 40*1800(T-288)

1233-T = 12(T-288)

1233-T = 12T - 3456

13T = 1233+3456

T = 360.69 K = 360.69-273 K = 87.69 K